OpenStudy (anonymous):

How do I give a "parametrization for the curve": the ray with initial point (2,5) that passes through (-1,0)?

OpenStudy (anonymous):

Sorry, I was eating.. To parametrize a curve, you are looking for some functions that relate to y and x. y = f(t) x = g(t) Then you just give bounds on t. The easiest way to do this is to write your original function y = f(x), then let x = t. So you have x = t, y = f(t) as your parametrization.

OpenStudy (anonymous):

So I find the equation of this ray, then substitute x with t? Btw, does it make a difference whether it is a ray or a line?... A line will be able to extend on both sides, but a ray will only extend on one.

OpenStudy (anonymous):

Right, so that means that t is limited. \(2 \le t \lt \infty\)

OpenStudy (anonymous):

Wouldn't it be \[-\infty < t \le 2\] Since the initial point is (2,5)? Do you mind checking my work? I got \[y=\left(\begin{matrix}5 \\ 3\end{matrix}\right)t+\left(\begin{matrix}25 \\ 3\end{matrix}\right)\] x=t and \[-\infty < t \le 2\]